OpenStudy (anonymous):

solve by factoring 10t^2 – t − 2 = 0

OpenStudy (anonymous):

(2t-1)(5t+2)=0 Hope this helps.

OpenStudy (anonymous):

Try to work it backwards now. (2t - 1)(5t + 2) = 0 (2t * 5t) + (2t*2) + (5t*1) + (2*-1) = 0 This will give you the original equation.